The angle of elevation from the tip of the shadow of a 12-foot pole to the top of the pole is 60 deg. How far is it from the tip of the shadow to the top of the pole . Round to the nearest tenth of a foot
Answer by Alan3354(21581)

sin(60) = 12/x
x = 12/sin(60) = 12/0.86603
x = 13.856 feet
x = 13.9 feet (to the nearest 0.1 foot)